OpenStudy (anonymous):

PLEASE HELP ME ON THIS MATH QUESTION :( :) Which of these has the greatest area? Use three point one four for pi. rhombus with side lengths: ten inches, height: 4 inches triangle with base length: twelve inches, height: 6 inches circle with radius: 4 inches trapezoid with base lengths: 3 inches and eleven inches, height: 6 inches

OpenStudy (anonymous):

Is it the circle?

OpenStudy (anonymous):

It s true Modom ! Exelent !:)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

the circle is 50.25 the triangle is 36 the rhombus is 40 the trapezoid is 42

OpenStudy (anonymous):

Your welcome ! I can say why if you want !

OpenStudy (anonymous):

bbbbbbbbbbbbbbb is right !

OpenStudy (anonymous):

Lol ok .... Thanks ... can you guys help me with another questions please?

OpenStudy (anonymous):

circle= π * r2 rhombus=base * height triangle= ½ * b * h trapeziod= ½(a+b) * h

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

OK hold on :)

OpenStudy (anonymous):

Say next questions !